### v1

# Low Photon Number Non-Invasive Imaging Through Time-Varying Diffusers

Abstract
Scattering poses a significant challenge in optical imaging. In microscopy, it leads to progressive degradation of image quality preventing sample examination at increasing depths. As the thickness of the sample increases, the number of photons that can successfully pass through it decreases, leading to a reduced signal and preventing us from obtaining clear images. Imaging becomes even more difficult with dynamic scatterers like biological tissues that change over time. We demonstrate non-invasive imaging of fluorescent objects hidden behind dynamic diffusers. Our reconstruction uses a stack of images recorded for different diffuser realizations and works even when individual camera frames contain small photon numbers.
keywords:
Imaging in dynamic complex media, Single photon detection, Non-invasive imaging, Speckle pattern analysis
1 Introduction
Scattering plays a crucial role in determining the appearance of observed objects. When scattering occurs multiple times within a medium, such as biological tissues, fog, or clouds, it is referred to as a complex medium. Light propagation through complex media results in random interference patterns known as speckle patterns[1]. For a long time, light scattering in complex media was considered uncontrollable. This fundamental challenge of light scattering affects not only imaging [2], but also poses obstacles in communication, spectroscopy, and other optical and photonic technologies.
Imaging through complex media has long been seen as a formidable challenge. However, a groundbreaking experiment by Vellekoop and Mosk [3] demonstrated that by controlling wavefronts through spatial light modulators (SLMs), light can be focused through scattering media. Subsequent experiments by Popoff et al. made significant advancements by introducing a method to measure the transmission matrix of complex media, allowing for wavefront control using SLMs behind these media [4, 5].
These remarkable advancements led to a better understanding of complex media and paved the way for developing non-invasive imaging techniques through such media. Specifically, experiments [6, 7] utilized the phenomenon known as the memory effect. Within small regions, when an input wavefront reaching a complex medium is tilted within a certain angular range, the output wavefront is equally tilted, resulting in the translation of the speckle pattern behind the complex medium [8, 9]. In imaging, within a memory effect region, light emitted from each point source behind the complex medium creates an almost identical speckle pattern that is shifted according to the position of the source origin.
Methods presented in [6, 7, 10] utilize the memory effect to reconstruct objects hidden behind complex media by calculating the spatial autocorrelation of speckle images and employing iterative phase retrieval algorithms. An alternative group of methods relies on fluctuating, random speckle illumination and observing the sample’s response [11, 12]. Some of these methods use Non-negative matrix factorization (NMF) to disentangle the temporal and spatial information about the sample [13, 11, 14]. More recently, well-established reflection-matrix techniques were applied to incoherent fluorescence imaging [12].
While several techniques for imaging in complex media were developed and represent significant achievements, their widespread adoption remains limited. One limitation is that these techniques typically work with static diffusers. In live bioimaging, biological tissues change over time, necessitating methods that can handle dynamic scattering. Similarly, LIDAR must operate in fog, which changes over time. Some approaches to overcome this issue involve using high-speed digital micromirror devices (DMDs) [15, 16, 17, 18, 19] as spatial light modulators (SLMs) for fast transmission matrix measurement. However, all these methods require high signal levels, which hinder their practical applications. For rapidly changing diffusers, one must acquire light over a short period, resulting in low light intensities being captured. This creates a growing need for robust imaging techniques that work in low-light scenarios through dynamic complex media.
Here, we demonstrate a non-invasive technique for imaging behind dynamic scattering layers. This technique is capable of operating in the low-photon regime and with single photon cameras, with much less than one photon per pixel on average . We demonstrate our technique in a widefield fluorescence microscope to image astrocyte cells stained with the common dye Alexa Fluor 568 through a dynamically changing diffuser, recording less than one photon per camera pixel per frame. Our method does not require wavefront shaping, transmission matrix characterization, or light focusing through the complex medium.
2 Principles
Figure 1: Concept of our method. (a) Schematic of our experiment. We record a movie of a luminescent object behind a changing diffuser. Each frame is assumed to be captured fast enough to treat the diffuser as static during the exposure. (b) Each frame $I_{i}$ is a convolution of the object $O$ with the speckle PSF $S_{i}$ of the varying diffuser. (c) The spatial Fourier transform of (b). (d) Averaging the Fourier transform moduli of all frames approximates the modulus of the Fourier transform of the object image without scattering. (e) The phase of the complex Fourier transform, lost in this process, is retrieved using a phase retrieval algorithm [20].
Our method for imaging through a dynamic diffuser is based on techniques such as those presented in [6, 7, 10]. The method involves recording a movie of a luminescent object behind a changing diffuser (see Figure 1a). We assume that the frame duration is short enough for the diffuser to remain static during each frame. We demonstrate that our approach enables robust reconstruction even if individual frames do not have a sufficient detected photon number for object reconstruction.
When the medium changes during the integration time, the resulting speckle pattern blurs, leading to an averaged-out image that loses spatial information about the sample. To overcome this, rather than directly integrating the signal recorded by the camera, we divide the signal into short-exposure, low-photon frames. We compute the Fourier transform modulus for each frame and average the resulting matrices using root mean square. This method enables effective imaging in dynamic media, yielding improved results, particularly in low-light scenarios, compared to the techniques described by [7, 6, 10].
Since each frame $I_{i}$ captured by the camera is a convolution of the object $O$ with the speckle Point Spread Function (PSF) $S_{i}$ of the varying diffuser (as illustrated in 1b), and the object is within the memory effect range [8, 9] (meaning each point of the object has the same PSF $S_{i}$), we have:
$I_{i}=O*S_{i},$
where $i=1,2,3,\ldots,N$ is the frame index, and $N$ is the total number of frames.
This implies that
$\mathcal{F}\{I_{i}\}=\mathcal{F}\{O\}\mathcal{F}\{S_{i}\},$
where $\mathcal{F}\{I_{i}\}$, $\mathcal{F}\{O\}$, and $\mathcal{F}\{S_{i}\}$ denote the spatial Fourier transforms of each frame $I_{i}$, the object $O$, and the speckle Point Spread Function $S_{i}$, respectively (as shown in 1c). Then,
$\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{I_{i}\}\right|^{2}}=\left|\mathcal{F}\{O\}\right|\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{S_{i}\}\right|^{2}}.$
For many realizations of the diffuser, the root mean square of the Fourier transform moduli of the frames for the object recorded through the diffuser converges to the modulus of the Fourier transform of the object image without scattering:
$\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{S_{i}\}\right|^{2}}\rightarrow\left|\mathcal{F}\{P\}\right|,$
where $\mathcal{F}\{P\}$ denotes the spatial Fourier transform of the Point Spread Function not affected by the diffuser (as demonstrated in Figure 1d). Note that this is different from direct imaging, because the NA is defined by the combination of the scatterer and the objective [21].
As a result of this analysis, the only information lost is the complex phase of the Fourier transform of the object image without scattering. As demonstrated previously [7, 6, 10], this phase can be retrieved using a phase retrieval algorithm (like in Figure 1e). In our work, we utilized the combination of hybrid input-output (HIO) and the error-reduction (ER) algorithms [20].
In practice, we didn’t directly average all the moduli of the Fourier transforms. Instead, we utilized several filtering techniques, such as removing the Gaussian envelope of the image and applying Hann windowing like in [7, 10]. For more details, please refer to the Methods section 5 and Supplemental Material.
3 Results
Figure 2: Imaging in the photon counting regime – simulation results. (a) Original object. (b) Image of the object captured with a finite resolution imaging system. (c) Modulus of the Fourier transform of the object (b). (d) Image of the object obscured by the diffuser - a single frame captured by the camera. (e) Single camera frame with Poisson noise. (f) An average of Fourier transforms modulus of 10,000 frames. (g) Reconstructed object image through a dynamic diffuser.
To demonstrate the effectiveness of our method in the single photon counting regime, we first conducted a computer simulation. Figure 2 presents the simulation results of imaging the object through the dynamic diffuser in a low photon number regime. Figure 2a shows an object consisting of several point emitters. Imaging this object through an imaging system with finite resolution gives the image shown in Figure 2b ($I=O*PSF$), with the Fourier transform modulus shown in Figure 2c (|$\left|\mathcal{F}\{I\}\right|=\left|\mathcal{F}\{O\}\right|\left|\mathcal{F}\{PSF\}\right|$).
Noiseless imaging of the object from Figure 2a through a diffuser results in the speckle image shown in Figure 2d ($I_{i}=O*S_{i}$). However, since each frame is short enough to capture the unchanged diffuser during its duration, the detected photon number is low. To account for this, we limit the number of photons detected during each frame from Figure 2d (adding a Poisson noise). Figure 2e shows a simulation of a single frame from Figure 2d with a limited number of photons.
In the simulation, we generated 10,000 frames, each with a different diffuser realization and an average of 400 photons per frame (0.04 photons per pixel). Figure 2f shows the root mean square of the Fourier transform moduli of these 10,000 frames ($\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{I_{i}\}\right|^{2}}=\left|\mathcal{F}\{O\}\right|\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left|\mathcal{F}\{S_{i}\}\right|^{2}}$). Notably, Figure 2f closely resembles Figure 2c. After applying the phase retrieval algorithm [20], we obtain the reconstructed image (Figure 2g), which closely resembles the direct image of the object in Figure 2b.
Figure 3: Imaging of Alexa Fluor 568 stained astrocytes through the dynamic complex medium. (a) Schematic of the wide-field fluorescence microscope utilized in our experiments. (b) Direct image of the sample without the diffuser (ground truth). (c) Single frame captured by the qCMOS camera. (d) Average of all 1872 frames recorded. (e) Reconstruction result using the phase retrieval algorithm [20].
In our experiment, we applied our method to fluorescence microscopy. As an experimental setup, we utilized a wide-field custom-made microscope (see Figure 3a). The details of the microscope construction can be found in Section 5. Our microscope allowed for the insertion and removal of a dynamic diffuser, which was placed just before the microscope objective. As a sensor in our microscope, we used the Hamamatsu ORCA-Quest 2 qCMOS camera, which enables single photon counting. We demonstrated our experimental setup on a sample of astrocytes stained with the popular fluorescent dye Alexa Fluor 568. Details of the sample preparation procedure are provided in Methods section 5. Figure 3b shows a direct image of the sample without the diffuser. Figure 3c illustrates a single frame captured by the qCMOS camera. During a frame duration of 200.96 ms, we recorded an average of 1.85 photons per camera pixel. Figure 3d shows the root mean square of all 1872 frames recorded in our experiment. Figure 3d corresponds to the incoherent sums of all frames. Speckles are blurred out and don’t exhibit any spatial information that could be retrieved. Therefore, we average the Fourier transform moduli of all the frames and apply a phase retrieval algorithm [20]. The reconstruction result using the phase retrieval algorithm [20] is presented in Figure 3e.
Figure 4: Imaging of fluorescence micro-spheres through the dynamic complex medium. (a) Direct images of fluorescent microspheres without the diffuser. (b) The modulus of the Fourier transforms of the images in column (a). (c) The average of the Fourier transform moduli of the frames captured through the dynamic diffuser. (d) Recovered image computed using the phase retrieval algorithm [20].
To demonstrate the versatility of our method, we applied it to various objects constructed with fluorescent micro-spheres samples (see Figure 4). The sample preparation procedure for the micro-spheres is detailed in Section 5. The column a of Figure 4 shows direct images of the fluorescent micro-spheres in our microscope. Figure 4b presents the modulus of the Fourier transforms of the images in the column a. Figure 4c shows the average of the Fourier transform modulus of the frames captured by the camera through the dynamic diffuser. Figure 4d displays the recovered image imaged through the diffuser, computed using the column c from fig. 4 with the phase retrieval algorithm [20].
4 Discussion
We successfully demonstrated imaging through a dynamic diffuser at signal levels low enough to detect single photons. Our method, like techniques presented in [7, 6, 10], does not require physical alteration of the sample, nor does it necessitate characterizing the medium’s transmission matrix or focusing light through it. However, our approach is specifically optimized for low-photon conditions and dynamic diffusers – a scenario highly relevant for deep tissue imaging. This approach can be implemented without adaptive optics or wavefront shaping systems.
We tested our method using a standard fluorescence microscope and applied it to various fluorescent samples. Notably, we successfully imaged astrocyte cells stained with the widely used dye Alexa Fluor 568, as well as various challenging objects constructed from fluorescent micro-spheres, further underscoring the versatility and effectiveness of our approach. Although we demonstrated our technique in microscopy, it has potential applications in various other fields of imaging. For instance, it can be utilized in LIDAR to cope with time-varying fog [22, 23], in astronomy for imaging distant, dim objects through a fluctuating atmosphere [24], and in endoscopy using multimode fibers [25, 26], where bends generate changing scattering scenarios.
Single camera frames in our experiment contain low signals, reaching $0.33$ photons per pixel per frame (around $10^{6}$ per frame). These experiments are feasible thanks to recent advances in the development of single photon cameras. The camera that we utilized has a relatively low frame rate, capturing fewer than 5 frames per second for the full field of view required by the speckle size. A potential improvement in the data acquisition rate our setup is the use of SPAD arrays, which are increasingly being applied in imaging [27, 28, 29, 30, 31]. Recent advancements in this field have led to the development of megapixel SPAD arrays [29], making their practical application in widefield microscopy feasible. Cameras based on SPAD technology allow for single photon detection with high temporal resolution, providing a powerful tool for enhancing various bioimaging techniques and other applications. In our application, orders of magnitude improvement of temporal resolution offered by SPAD arrays will enable imaging through extremely fast varying diffusers.
5 Methods
5.1 Experimental setup
We designed and constructed our experimental setup, which consisted of a fluorescence wide-field microscope. For the objective, we used Nikon S Plan Fluor ELWD 20x/0.45, and for the excitation beam, we utilized a Spectra-Physics Millennia PRO laser with a wavelength of 532 nm. We utilized several neutral density filters to attenuate the laser beam. The fluorescence sample was placed within the working distance of the primary objective, while a rotating diffuser (Edmund Holographic Diffuser of 5° Diffusing Angle) was positioned adjacent to the microscope objective. To select the desired field of view in our experiment, we illuminated the sample from a transmissive side using another objective (Nikon Nikon S Plan Fluor ELWD 40x/0.60), which was usually out of focus to provide a large excitation spot.
To filter the fluorescence light from the excitation source, we employed a dichroic mirror (Edmund 552), a long-pass filter (Thorlabs FELH0550), and a notch filter Semrock NF01-532U-25. Additionally, to ensure good visibility of speckles, a band-pass filter (Thorlabs FBH590-10) was used. The tube lens in our setup was a Thorlabs TTL200-A with a focal length of 200 mm. The ORCA-Quest 2 qCMOS camera from Hamamatsu was positioned at the focal point of the tube lens.
Mounts for the fluorescence sample and the rotating diffuser with its motor were custom-designed and 3D-printed.
For more details of the experimental setup, please refer to Supplementary information.
5.2 Data acquisition
In our experiment, we utilized a galvo mirror (Thorlabs GVS211/M) positioned in the excitation laser beam path to modulate the illumination of the sample. The galvo mirror was connected to a microcontroller (Arduino Micro), which communicated with a computer via a serial port. A Python program controlled both the microcontroller and the camera.
Initially, the computer sent a command to the Arduino to illuminate the sample and then to the camera to begin capturing light. The microcontroller ensured that the fluorescent sample was illuminated only during the camera’s light detection period to limit the photo-bleaching of the sample. It was removing the excitation laser beam immediately after the camera sent a trigger indicating the end of the frame, ensuring precise synchronization.
The microcontroller used in our experiment had an additional task: controlling the rotation of the diffuser. The microcontroller was connected to an electronic circuit (L2930) that enabled it to switch the 12V power supply required for motor (SKU GS06331-10) rotation on and off.
5.3 Data analysis
We performed several filtering techniques in our data processing to retrieve the object.
First, we divided by a 2D Gaussian function. The speckle image on the camera is not equally distributed across the sample, exhibiting a Gaussian envelope. By dividing by a 2D Gaussian function, we corrected for this uneven distribution.
Next, we applied a 2D Hann filter to smooth the edges of each frame, which helps mitigate edge effects in the Fourier domain.
After applying the Hann filter, we computed each frame’s 2D Fourier transform and calculated its modulus. We then averaged with the root mean square these transformed frames and convolved the resulting signal with a narrow 2D filter to remove speckle artifacts.
Finally, we applied the phase retrieval algorithm (the combination of hybrid input-output (HIO) and the error-reduction (ER) algorithms) to the prepared Fourier amplitude, to obtain a reconstructed object.
5.4 Sample preparation
Fluorescence micro-spheres
Fluorescent microsphere samples were prepared using a drop-casting method. A concentrated solution of fluorescent microspheres (PS-FluoRed-Particles; 4.99 µm/SD=0.16 µm; 530/607 nm) was significantly diluted with ethanol. The solution was subjected to a 15-minute sonication bath to prevent the clustering of the fluorescent microspheres. A 50 µL drop of this diluted solution was then placed onto a microscope coverslip and allowed to air dry.
Alexa Fluor 568 stained astrocyte
Three-week-old primary mixed neuronal cultures on glass slides were fixed in PBS with 4% paraformaldehyde and 4% sucrose for 8 minutes at $37^{o}$. After washing three times with PBS containing 4% sucrose, cells were permeabilized with 0.25% Triton X-100 for 2 minutes. Blocking was performed in PBS with 3% BSA and 0.1% Triton X-100 for 1 hour at room temperature. Cultures were incubated overnight at $4^{o}$ with an anti-GFAP primary antibody (1:1000, #ab7260), followed by three PBS washes. Secondary antibody incubation was conducted for 1 hour at room temperature using Alexa Fluor 568-conjugated anti-rabbit antibody (1:500, #A-11011). After three PBS washes, samples were mounted on glass slides using Fluoromount-G solution (#00-4958-02).
[…]
Acknowledgements We would like to express our sincere gratitude to Diana Legutko for preparing the Alexa Fluor 568-stained astrocyte samples, and to Anat Daniel for insightful discussions and comments about the experimental setup. We also acknowledge the support of the following funding agencies: Narodowe Centrum Nauki (grants 2023/49/N/ST7/04195 and 2022/47/B/ST7/03465); Fundacja na rzecz Nauki Polskiej (FIRST TEAM project POIR.04.04.00-00-3004/17-00); the European Regional Development Fund (POIR.04.04.00-00-3004/17-00); HORIZON EUROPE Marie Skłodowska-Curie Actions (FLORIN ID 101086142); and the French Government Scholarship for Ph.D. Cotutelle/Codirection.