### v2

# Low Photon Number Non-Invasive Imaging Through Time-Varying Diffusers

Abstract
Optical imaging plays a crucial role in advancing science and technology, enabling applications in fields ranging from biomedicine to astronomy. However, imaging through scattering media such as biological tissues, fog, or turbulent atmosphere remains a major challenge. Light scattering and absorption in such media make imaging challenging; in the case of time varying scatterers and low-light regime imaging has not been demonstrated so far. We present the first demonstration of non-invasive imaging of dim objects hidden behind dynamic scattering layers, obtaining robust reconstruction even at extremely low photon counts per frame. We achieve this by developing a new data-processing approach. In our experiment, we utilize a photon number resolving camera to capture a sequence of frames, containing on average, less than one photon per pixel. We validate our approach in microscopy, where we reconstruct images of biological samples stained with standard fluorescent dyes. Beyond microscopy, our approach can be applied in LIDAR systems for imaging through fog, and endoscopy using multimode and multicore fibers.
1 Introduction
Light scattering plays a crucial role in determining the appearance of most of the things we see. When scattering occurs multiple times within a medium, such as biological tissues, fog, or clouds, it is called a complex medium. The propagation of light through complex media results in random interference patterns known as speckle patterns[1]. For a long time, light scattering in complex media was considered uncontrollable. This fundamental challenge of light scattering affects not only imaging [2], but also poses obstacles in communication, spectroscopy, and other optical and photonic technologies.
Imaging through complex media has long been seen as a formidable challenge. However, a groundbreaking experiment by Vellekoop and Mosk [3] demonstrated that by controlling wavefronts through spatial light modulators (SLMs), light can be focused through scattering media. Subsequent experiments by Popoff et al. significantly advanced the field by introducing a method to measure the transmission matrix of complex media [4, 5]. This development enabled light control behind those media using SLMs.
These remarkable advancements led to a better understanding of complex media and paved the way for developing non-invasive imaging techniques through such media. Specifically, experiments by Bertolotti et al. or Katz et al. [6, 7] utilized the phenomenon known as the memory effect. Within small regions, when an input wavefront reaching a complex medium is tilted within a certain angular range, the output wavefront is equally tilted, resulting in the translation of the speckle pattern behind the complex medium [8, 9]. In imaging, within a memory effect region, light emitted from each point source behind the complex medium creates an almost identical speckle pattern that is shifted according to the position of the source origin.
Methods presented in [6, 7, 10] utilize the memory effect to reconstruct objects hidden behind complex media by calculating the spatial autocorrelation of speckle images and employing iterative phase retrieval algorithms. An alternative group of methods relies on fluctuating, random speckle illumination and observing the sample’s response [11, 12]. Some of these methods use Non-negative matrix factorization (NMF) to disentangle the temporal and spatial information about the sample [13, 11, 14]. More recently, well-established reflection-matrix techniques were applied to incoherent fluorescence imaging [12].
While several techniques for optical imaging in complex media were developed and represent significant achievements, their widespread adoption remains limited. One limitation is that these techniques typically work with static diffusers. In live bioimaging, biological tissues change over time, necessitating methods that can handle dynamic scattering. Similarly, LIDAR must operate in fog, where light scattering changes over time. Some approaches to overcome this issue involve using high-speed digital micromirror devices (DMDs) [15, 16, 17, 18, 19] as SLMs for fast transmission matrix measurement. However, all these methods require high signal levels, which hinder their practical applications. For rapidly changing diffusers, one must acquire light over a short period, to approximate the diffuser as static. This inevitably results in low light intensities being captured. This creates a growing need for robust imaging techniques that work in low-light intensity scenarios through dynamic complex media.
We demonstrate that non-invasive imaging behind dynamic scattering layers is possible even in the low-photon number regime. There was no previous demonstration under such conditions. To achieve this, we develop a data-analysis approach that reconstructs the object from a sequence of ultra-dim frames recorded behind a scattering layer, without ballistic light. Our imaging method builds on concepts introduced in earlier works [6, 7, 10], but extends their capabilities to enable imaging through dynamically changing diffusers and under low-light conditions.
The proposed approach operates even under extreme cases, only two photons per frame, and remains effective in highly dynamic scenarios, as it does not require the scattering medium to remain static between frames. We present a proof-of-concept experimental demonstration of our method in a widefield microscope imaging fluorescent samples. We use a single-photon camera operating in an ultra-low-light regime, recording fewer than one photon per pixel per frame. Our method does not require wavefront shaping, transmission matrix characterization, or light focusing through the complex medium.
2 Principles
Figure 1: Concept of the method. (a) A luminescent, spatially incoherent object is imaged through a time-varying diffuser and recorded with a high-speed, photon-counting camera. (b) A long exposure produces a homogeneous image that carries no spatial information about the object. (c) A single short-exposure frame preserves speckle structure but contains very few photons. (d) In the analysis, we use a series of short frames with a low signal. (e) Summing many such low-signal frames similarly yields a homogeneous image without spatial content. (f) Instead, for each short-exposure frame, we compute the magnitude of its Fourier transform. (g) The root-mean-square (RMS) average of these Fourier magnitudes across the burst recovers the spatial-frequency content of the object. (h) After retrieving the missing Fourier phase using a phase-retrieval algorithm [20], the object hidden behind the dynamic diffuser can be reconstructed.
Our approach involves recording a movie of a luminescent object placed behind a changing diffuser (see Fig. 1a). When the camera integration time is long, the diffuser undergoes substantial changes during a single exposure (see Fig. 1b). As a result, the recorded speckle patterns average out, yielding a spatially homogeneous image that carries no information about the hidden object. If the exposure time is short, the frame can be approximated as corresponding to a single realization of the diffuser. However, such frames may contain very low photon counts, sometimes down to a few detected photons per frame (see Fig. 1c).
In our method, we acquire a sequence of short exposure frames (see Fig. 1d). A naive integration of these frames would produce an image similar to that obtained with a single long-exposure measurement (see Fig. 1e), again lacking spatial information. Instead, for each frame we compute the modulus of its two-dimensional Fourier transform (see Fig. 1f). We then average these Fourier-modulus images (see Fig. 1g). Remarkably, the resulting averaged matrix corresponds to the Fourier-transform modulus of the object hidden behind the diffuser and can be obtained even when individual frames contain only single detected photons. With this modulus, the object image can be reconstructed (see Fig. 1h) using well established phase-retrieval algorithms.
This method enables effective imaging under dynamic scattering media conditions and achieves reliable results in low-photon regimes that allow tackling extreme low-light and dynamic scattering conditions, which are completely beyond the reach of currently demonstrated techniques.
Figure 2: Mathematical background of our method. (a) Direct imaging in a scattering-free system. The image $I$ formed on the camera is a convolution of the object $O$ with the system Point Spread Function (PSF) $h$. (b) Under dynamic scattering, each speckle image $I_{i}$ formed on the camera is a convolution of the object $O$ with the instantaneous speckle Point Spread Function (PSF) $S_{i}$. (c) Spatial Fourier transforms of the images shown in (b). (d) The root mean square (RMS) of the Fourier transform moduli, computed across many frames, converges to $|\mathcal{F}\{O\}|\,|\mathcal{F}\{P\}|$, where $P$ denotes the effective PSF of the system as if no scattering were present. (e) The magnitude of the scattering-free Fourier transform is obtained, while its phase is lost. A phase retrieval algorithm [20] is then used to recover the missing phase information. (f) Reconstruction of the object after phase retrieval.
In a conventional, scattering-free imaging system, the image recorded by the camera $I$ is a 2D mathematical convolution of the object $O$ and the system’s Point Spread Function (PSF) $h$ (see Fig. 2a). In the presence of scattering, however, the situation becomes more complex. The pattern formed on the camera, known as a speckle pattern, does not resemble an ordinary PSF. Nevertheless, for small objects located within the memory-effect range [8, 9], each point of the object produces the same speckle pattern, shifted relative to one another. As a result, the image captured by the camera is the convolution of the object with this speckle pattern, which acts as a new speckle Point Spread Function.
Under dynamic scattering conditions, each speckle image formed on camera is a convolution of the object $O$ with the instantaneous speckle Point Spread Function (PSF) $S_{i}$ of the varying diffuser (as illustrated in Figure 2b). We can represent each speckle image as:
$I_{i}=O*S_{i},$ (1)
where $i=1,2,3,\ldots,M$ is the image index, and $M$ is the total number of images. Note that, each frame $F_{i}$ captured by the camera is a Poisson-distributed measurement of the underlying noiseless speckle image.
Under this model, the Fourier transforms satisfy:
$\mathcal{F}\{I_{i}\}=\mathcal{F}\{O\}\mathcal{F}\{S_{i}\},$ (2)
where $\mathcal{F}\{I_{i}\}$, $\mathcal{F}\{O\}$, and $\mathcal{F}\{S_{i}\}$ denote the spatial Fourier transforms of each image $I_{i}$, the object $O$, and the instantaneous speckle Point Spread Function $S_{i}$, respectively (as shown in Fig. 2c). By computing the root mean square of the Fourier transform moduli across all images (Fig. 2d), we obtain:
$\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{I_{i}\}\right|^{2}}=\left|\mathcal{F}\{O\}\right|\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{S_{i}\}\right|^{2}}.$ (3)
For many realizations of the diffuser, the root mean square of the Fourier transform moduli of speckle Point Spread Functions $S_{i}$ averages [1] to Fourier transform of the effective Point Spread Function ($\mathcal{F}\{P\}$) of the imaging system as it would not have been affected by the diffuser (see Fig. 2d). Among common averaging strategies, the RMS operator provides the fastest and most reliable convergence to the true scattering-free Fourier magnitude, as demonstrated by our statistical analysis in the Supplementary Information, Section 2.1.2.
$\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{S_{i}\}\right|^{2}}\;\xrightarrow[M\to\infty]{}\;\left|\mathcal{F}\{P\}\right|.$ (4)
Because each recorded frame $F_{i}$ is a Poisson-distributed measurement of the underlying noiseless speckle image, its Fourier magnitude also satisfies a law-of-large-numbers convergence. Therefore,
$\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{F_{i}\}\right|^{2}}\;\xrightarrow[M\to\infty]{}\;\left|\mathcal{F}\{O\}\right|\,\left|\mathcal{F}\{P\}\right|,$ (5)
In the Supplementary Information Section 2.1.2, we provide a comprehensive statistical evaluation, estimating the required number of frames for image reconstruction, the achievable resolution, and the signal-to-noise ratio (SNR) of the recovered Fourier magnitude.
Note that in this setting, imaging through the diffuser is different from direct imaging, because the effective numerical aperture NA is defined by the combination of the scatterer and the objective [21].
As a result of this analysis, we retrieve the magnitude of the Fourier transform of the image of the object that is not affected by scattering (see Fig. 2e), and information about its complex phase is lost. As demonstrated previously [7, 6, 10], this phase can be retrieved using a phase retrieval algorithm. In our work, we utilized the combination of hybrid input-output (HIO) and the error-reduction (ER) algorithms [20]. Once the Fourier phases are estimated, the object obscured by the dynamic diffuser can be retrieved (like in Figure 2f).
We detail the post-processing techniques and algorithms used in our analysis in Supplementary Information Sections 2.1.4 and 3.
It is important to note that, even though phase retrieval algorithms are highly sensitive to noise, we do not apply them to noisy raw frames. Instead, we apply phase retrieval to the estimated high-SNR Fourier modulus of the object hidden behind a dynamically changing scatterer. For more details of data analysis, please refer to the Methods Section 5 and Supplementary Information Section 3.
3 Results
Figure 3: Imaging in the photon counting regime – simulation results. (a) Original object. (b) Image of the object captured with a finite resolution imaging system. (c) Modulus of the Fourier transform of the object (b). (d) Image of the object obscured by the diffuser - a single frame captured by the camera. (e) Single camera frame with Poisson noise. (f) An average of Fourier transforms modulus of 10,000 frames. (g) Reconstructed object image through a dynamic diffuser.
To demonstrate the effectiveness of our method in the single photon counting regime, we first conducted a computer simulation. Figure 3 presents the simulation results of imaging the object through the dynamic diffuser in a low photon number regime. Figure 3a shows an object consisting of several point emitters. Imaging this object through an imaging system with finite resolution gives the image shown in Figure 3b (I=OPSFI=O*PSF), with the Fourier transform modulus shown in Figure 3c ($\left|\mathcal{F}\{I\}\right|=\left|\mathcal{F}\{O\}\right|\left|\mathcal{F}\{PSF\}\right|$).
Noiseless imaging of the object from Figure 3a through a diffuser results in the speckle image shown in Figure 3d ($I_{i}=O*S_{i}$). However, since each frame is short enough to capture the unchanged diffuser during its duration, the detected photon number is low. To account for this, we limit the number of photons detected during each frame from Figure 3d (adding a Poisson noise). Figure 3e shows a simulation of a single frame from Figure 3d with a limited number of photons.
In the simulation, we generated 10,000 frames, each with a different diffuser realization and an average of 400 photons per frame (0.04 photons per pixel). Figure 3f shows the root mean square of the Fourier transform moduli of these 10,000 frames ($\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{I_{i}\}\right|^{2}}=\left|\mathcal{F}\{O\}\right|\sqrt{\frac{1}{M}\sum_{i=1}^{M}\left|\mathcal{F}\{S_{i}\}\right|^{2}}$). Notably, Figure 3f shows excellent agreement with Figure 3c. After applying the phase retrieval algorithm [20], we obtain the reconstructed image (Figure 3g), which corresponds to the direct image of the object in Figure 3b.
[…]
In our experiment, we applied our method to fluorescence microscopy. As an experimental setup, we utilized a wide-field custom-made microscope (see Figure 4a). The details of the microscope construction can be found in Section 5. Our microscope allowed for the insertion and removal of a dynamic diffuser, placed just before the microscope objective (between the objective and the fluorescent sample). As a sensor in our microscope, we used the Hamamatsu ORCA-Quest 2 qCMOS camera, offering photon number resolution at each pixel.
[…]
Figure 5: Results comparison of imaging through the dynamic complex medium. (a) Direct images of fluorescent samples without the diffuser. (b) The modulus of the Fourier transforms of the images in column (a). (c) Example camera frames of objects in column (a) obscured by the varying diffuser, with photon counts as low as 0.14 photons per pixel in 1152 frames, highlighted in the second row. (d) The average of the Fourier transform moduli of the frames captured through the dynamic diffuser. (e) Recovered image computed using the phase retrieval algorithm [20]. Rows 1 to 3 present fluorescent microspheres, while row 4 shows Alexa Fluor 568-labeled astrocytes.
To demonstrate the versatility of our method, we applied it to various objects composed of fluorescent micro-spheres and astrocytes stained with the popular fluorescent dye Alexa Fluor 568 (see Figure 5). Fluorescent samples preparation procedures are detailed in Section 5.4. The column a of Figure 5 shows direct images of the fluorescent samples in our microscope. Figure 5b presents the modulus of the Fourier transforms of the images in the column a. Figure 5c presents individual frames for different objects captured through the varying diffuser, with photon counts reaching as low as 0.14 photons per pixel in 1152 captured frames, as shown in the second row. Figure 5d shows the average of the Fourier transform modulus of the frames captured by the camera through the dynamic diffuser. Figure 5e displays the recovered image imaged through the diffuser, computed using the column d from fig. 5 with the phase retrieval algorithm [20]. Rows 1 to 3 present results obtained with fluorescent microspheres, while row 4 shows data from Alexa Fluor 568-labeled astrocytes.
4 Discussion
We successfully demonstrated imaging through a dynamic diffuser at signal levels low enough to detect single photons, conditions under which no imaging has been demonstrated so far. This scenario is highly relevant for deep-tissue imaging, where scattering and low photon counts are significant challenges. Our approach does not require physical alteration of the sample, nor does it necessitate characterizing the medium’s transmission matrix or focusing light through it. Our approach does not require adaptive optics or wavefront shaping systems. Our method is robust, remaining unaffected by spatially uncorrelated noise.
Moreover, our method is resilient to fluorescence bleaching and minor sample movements during multi-frame acquisition. Although bleaching reduces signal intensity, the spatial distribution of fluorophores in densely stained samples remains intact, preserving essential spatial information despite reduced Fourier amplitudes. By averaging the Fourier transform moduli across frames, we accurately recover the object’s Fourier transform. Similarly, since translational motion only introduces phase in the Fourier domain, our modulus-based approach naturally compensates for such shifts—only significant rotations or structural changes would impact the reconstruction. Our method works also in the case of static diffusers (see Supplementary Information Section 2.1.6).
In Section 2 of the Supplementary Information, we provide a detailed discussion of the theoretical foundations of our approach. We analyze the required number of frames for a given photon count per frame and demonstrate that our approach has no strict lower bound on signal intensity, allowing object reconstruction by accumulating a sufficient number of frames. In addition, we explore the spatial resolution of our imaging technique and its signal-to-noise ratio.
We validated our approach using a standard fluorescence microscope and applied it to various fluorescent samples. Notably, we successfully imaged astrocyte cells stained with the widely used dye Alexa Fluor 568, as well as various challenging objects constructed from fluorescent micro-spheres, further underscoring the versatility and effectiveness of our approach. Our approach has potential applications in various other fields of imaging. For instance, it can be utilized in LIDAR to cope with time-varying fog [22, 23], in astronomy for imaging distant, dim objects through a fluctuating atmosphere [24], and in endoscopy using multimode fibers [25, 26], where bends generate changing scattering scenarios.
Our approach has no lower bound on fluorescence intensity and can operate even with just 2 photons per frame. However, achieving performance with as few as 2 photons per frame may require a substantial number of frames, as detailed in Section 2.1.6 of the Supplementary Information. Single camera frames in our experiment contain low signals, reaching $0.14$ photons per pixel per frame (around $3\times 10^{5}$ per 9 Mpx frame). These experiments are feasible thanks to recent advances in the development of single photon cameras. The camera that we utilized has a relatively low frame rate, capturing fewer than 5 frames per second for the full field of view required by the speckle size. Since our experiment requires thousands of camera frames, the total acquisition time for a single image reconstruction is typically a couple of minutes. Nonetheless, this acquisition duration is well-suited for many bioimaging applications, such as embryonic development, cell division, and synaptic modifications.
A potential improvement in the data acquisition rate our setup is the use of SPAD arrays, which are increasingly being applied in imaging [27, 28, 29, 30, 31]. Recent advancements in this field have led to the development of megapixel SPAD arrays [29], making their practical application in widefield microscopy feasible. Cameras based on SPAD technology allow for single photon detection with high temporal resolution, providing a powerful tool for enhancing various bioimaging techniques and other applications. Some SPAD arrays offer sub-picosecond resolution for photon arrival time tagging [29]. By correlating detected photons within a fixed time window while the diffuser remains stationary, we can optimize data utilization and significantly enhance acquisition efficiency, enabling much faster fluorescence imaging. In our case, orders of magnitude improvement of temporal resolution offered by SPAD arrays will enable imaging through extremely fast varying diffusers, making it suitable for more realistic scenarios.
5 Methods
5.1 Experimental setup
We designed and constructed our experimental setup, which consisted of a fluorescence wide-field microscope. For the objective, we used Nikon S Plan Fluor ELWD 20x/0.45, and for the excitation beam, we utilized a Spectra-Physics Millennia PRO laser with a wavelength of 532 nm. We utilized several neutral density filters to attenuate the laser beam. The fluorescence sample was placed within the working distance of the primary objective, while a rotating diffuser (Edmund Holographic Diffuser of 5° Diffusing Angle) was positioned adjacent to the microscope objective (between the objective and the sample). To select the desired field of view in our experiment, we illuminated the sample from a transmissive side using another objective (Nikon S Plan Fluor ELWD 40x/0.60), which was usually out of focus to provide an adjustable excitation spot.
To separate the fluorescence light from the excitation source, we employed a dichroic mirror (Edmund 552), a long-pass filter (Thorlabs FELH0550), and a notch filter Semrock NF01-532U-25. Additionally, to ensure good visibility of speckles, a band-pass filter (Thorlabs FBH590-10) was used. The tube lens in our setup was a Thorlabs TTL200-A with a focal length of 200 mm. The ORCA-Quest 2 qCMOS camera from Hamamatsu was positioned at the focal point of the tube lens.
Mounts for the fluorescence sample and the rotating diffuser with its motor were custom-designed and 3D-printed.
For more details of the experimental setup, please refer to the Supplementary Information Section 1.
5.2 Data acquisition
In our experiment, we utilized a galvo mirror (Thorlabs GVS211/M) positioned in the excitation laser beam path to modulate the illumination of the sample. The galvo mirror was connected to a microcontroller (Arduino Micro), which communicated with a computer via a serial port. A Python program controlled both the microcontroller and the camera.
Initially, the computer sent a command to the Arduino to illuminate the sample and then to the camera to begin the acquisition. The microcontroller ensured that the fluorescent sample was illuminated only during the camera’s light detection period to limit the photo-bleaching of the sample. It was removing the excitation laser beam immediately after the camera sent a trigger indicating the end of the frame.
The microcontroller used in our experiment had an additional task: controlling the rotation of the diffuser. The microcontroller was connected to an electronic circuit (L2930) that enabled it to switch the 12V power supply required for motor (SKU GS06331-10) rotation on and off.
5.3 Data analysis
We performed several filtering techniques in our data processing to retrieve the object. The speckle image on the camera is not equally distributed across the sample, exhibiting a Gaussian envelope. By dividing by a 2D Gaussian function, we corrected for this uneven distribution. Next, we applied a 2D Hann filter to smooth the edges of each frame, which helps mitigate edge effects in the Fourier domain. After applying the Hann filter, we computed each frame’s 2D Fourier transform and calculated its modulus. We then averaged these transformed frames with the root mean square and convolved the resulting signal with a narrow 2D filter to remove speckle artifacts. Finally, we applied the phase retrieval algorithm (the combination of hybrid input-output (HIO) and the error-reduction (ER) algorithms) to the prepared Fourier amplitude, to obtain a reconstructed object.
5.4 Sample preparation
Fluorescence micro-spheres
Fluorescent microsphere samples were prepared using a drop-casting method. A concentrated solution of fluorescent microspheres (PS-FluoRed-Particles; 4.99 µm/SD=0.16 µm; 530/607 nm) was significantly diluted with ethanol. The solution was subjected to a 15-minute sonication bath to prevent clustering of the fluorescent microspheres. A 50 µL drop of this diluted solution was then placed onto a microscope coverslip and allowed to air dry.
Alexa Fluor 568 stained astrocyte
Three-week-old primary mixed neuronal cultures on glass slides were fixed in PBS with 4% paraformaldehyde and 4% sucrose for 8 minutes at $37^{o}$C. After washing three times with PBS containing 4% sucrose, cells were permeabilized with 0.25% Triton X-100 for 2 minutes. Blocking was performed in PBS with 3% BSA and 0.1% Triton X-100 for 1 hour at room temperature. Cultures were incubated overnight at $4^{o}$C with an anti-GFAP primary antibody (1:1000, #ab7260), followed by three PBS washes. Secondary antibody incubation was conducted for 1 hour at room temperature using Alexa Fluor 568-conjugated anti-rabbit antibody (1:500, #A-11011). After three PBS washes, samples were mounted on glass slides using Fluoromount-G solution (#00-4958-02).
[…]
Data availability The data used in this study is available from the corresponding author upon reasonable request.
[…]
Acknowledgements We would like to express our sincere gratitude to Diana Legutko for preparing the Alexa Fluor 568-stained astrocyte samples, and to Anat Daniel for insightful discussions and comments about the experimental setup. We want to acknowledge the support of the following funding agencies: Foundation for Polish Science (FIRST TEAM project FENG.02.02-IP.05-0253/23); Narodowe Centrum Nauki (grants 2023/49/N/ST7/04195 and 2022/47/B/ST7/03465); National Centre for Research and Development QuantERA II projects: QM3 (QuantERAII/02/QM3/03/2024) and EXTRASENS (QuantERAII/02/EXTRASENS/02/2024); HORIZON EUROPE Marie Skłodowska-Curie Actions (FLORIN ID 101086142); and the French Government Scholarship for Ph.D. Cotutelle/Codirection.‘
[…]
Conflict of interest The authors declare no competing interests.
[…]
Contributions A.M. contributed to conceiving the project idea together with S.G. and R.L., designed and built the experimental setup, prepared the fluorescence microsphere samples, conducted the experiments, performed all simulations, developed and investigated the data analysis methods, analyzed the data, prepared all figures, and wrote the manuscript. W.Z. contributed to the theoretical framework, supported improvements in the data analysis, and authored the supplementary section on theoretical investigations. P.S. provided key input on constructing the experimental platform, contributed to data acquisition and theory, and wrote the supplementary section detailing the experimental setup. B.G. supported the theoretical investigations and co-authored relevant parts of the theoretical supplement. S.G. contributed to its initial conception and provided critical insights and guidance throughout the project. R.L. conceived the project, coordinated its development, and provided essential guidance and input across all stages. All authors reviewed and revised the manuscript and supplementary materials.